OpenStudy (anonymous):

Write an equation in slope-intercept form for the line through the given points. Be sure to show all the work! A (4, 2) B (6, -3)

OpenStudy (gorv):

y=mx+c standard equation of line m= slope

OpenStudy (gorv):

(x1,y1)=(4,2) (x2,y2)=(6,-3)

OpenStudy (gorv):

slope=(y2-y1)/(x2-x1)

OpenStudy (gorv):

and put any point in eauation to find the value of c ...after calculating slope